Intellus has long-term debt of $5 million, owners' equity of $7.5 million, current assets of $1 million, gross fixed assets of $20 million, and accumulated depreciation of $7 million. What is the firm’s net working capital? Intellus has long-term debt of $5 million, owners' equity of $7.5 million, current assets of $1 million, gross fixed assets of $20 million, and accumulated depreciation of $7 million. What is the firm’s net working capital?

show formulas work Please! Thank You!
Answer #1

Calculation of current liabilities:

Current liabilities=Total assets- long term debt-owners equity

Current liabilities= (1+20-7)-5-7.5= 1.5 million

Calculation of working capital

Working capital= current assets- current liabilities= 1-1.50= -0.5 million

Working capital= -$0.50 million